colette has some 8% acid to which she adds 10 liters of 20% acid to get a 15% acid solution determine how much 8% acid she needs.
Answer by Paul(988)

8x+20(10)=15(x+10)
8x+200=15x+150
-7x=-50
x=7.14 (rounded)

Hence, she needs about 7.14L of 8% solution.
Paul.